5-41: Stock Rights. Martha Lou owns 100 shares of Blain Corporation common stock. She purchased the stock on July 25, 1998, for $4,000. On May 2 of the current year, she receives a nontaxable distribution of 100 stock rights. Each stock right has a $10 FMV, and the FMV of the Blain common stock is $70 per share. With each stock right, Martha Lou may acquire one share of Blain common for $68 per share. Assuming that she elects to allocate basis to the stock rights, answer the following:

a. What is the basis allocated to the stock rights?

b. If she sells the stock rights on June 10 for $1,080, determine the amount and character of the recognized gain?

c. If she exercises the stock rights on May 14, what is the basis of the 100 shares purchased and when does the holding period start?

d. If she does not elect to allocate basis to the stock rights, determine the amount and character of the gain if she sells the stock rights on June 10 for $1,080?
Answer #1

a)

The fair market value of each stock right is $10.

FMV of each common stock is $70.

The no. of shares is 100

Basis of stock right received

= ((10 *100) / (10+70)*100) * Purchase cost

= ((10*100) / (10+70)*100)*4000

= $500

b)

Capital Gain = amount realized - basis

= 1080-500

= $580

c)

Basis = Basis from (part a) + Purchase cost

= 500+68*100

= $7300

d)

Capital Gain = Amount realized - Basis

= 1080-0

= $1080

Therefor, Martha enjoys a LTCG of $1,080.